1. Complete each equation with a number so that the equation has the stated solution(s).
a) 𝑥2+7=𝑥2+_____; no solution
b) 𝑥2+7=𝑥2+_____ ; all real numbers
c) 𝑥2+7=_____;𝑥=±3

Answer by ikleyn(52786)

1. Complete each equation with a number so that the equation has the stated solution(s).
a) x2+7 = x2+_____; no solution
b) x2+7 = x2+_____ ; all real numbers
c) x2+7 =_____; x = ±3
~~~~~~~~~~~~~~~~~~~~


(a)  x^2 + 7 = x^2 + 8      (no solution)


(b)  x^2 + 7 = x^2 + 7      (all real numbers are solutions)


(c)  x^2 + 7 = 16           (the solutions are x = ±3)

I answered all your questions.